OpenStudy (anonymous):

given that angle A are SUPPLEMENTARY if angle A = (x+28) and angle B =(3x) Find angle A

OpenStudy (anonymous):

x= 38 so, A = x+28 = 66 and B = 3x = 114 A+B=114+66 = 180 (they are supplementary)